

********************************************************************
UARS WEEKLY STATUS REPORT - 23 FEB 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS was activated in its full power mode on January 11. It was
change to half power mode on January 28 when the AXIS-2 unit was deactivated.
AXIS-1 was again active in the half power mode between January 28 and
February 7. At the present time, energy deposition from PEM/AXIS is not being
computed due to a software problem which PEM is to repair.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, the source of these failures has been
traced to quota limits and these have been corrected. PEM Telemetry is runing
in duplicate streams through PEM:: and PEM2K::. We find that files which
transfer from the CDHF to PEM:: and PEM2K:: are now identical. PEM has
requested DAP command capability with the CMS from PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The instrument operations schedule for the next yaw period,
15 March to 19 April, has been distributed and posted on the
web page http://umpgal.gsfc.nasa.gov (Instrument Schedule button).
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 February 2001 (Orbit 51481, GMT Day 041) through 16
February 2001 (Orbit 51585, GMT Day 040).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 042/0134 51496 Power ON
SUSIM 042/0234 51497 Power ON, Started
PEM 042/1154 51503 ZEPS/HEPS Day-Only
SUSIM 042/1622 51506 Power ON Complete
AXIS 043/0053 51511 Power ON, 7w
HALOE 045/0001 51540 Power ON, part 1
HALOE 045/0429 51543 Power ON, part 2
PEM/AXIS 047/2320 51585 AXIS Power OFF
The beta angle ranged between the angles of 02.1 degrees to 25.8 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 48 degrees on 26 February 2001 (DOY 057).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 and 1.3 degrees C. Additional battery performance data for each
battery on 09 February 2001 and 16 February 2001 are provided in Appendix D.
The clock error ranged from ~9.8 to 10.2 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 10
¥ HALOE = 02
¥ SUSIM = 02
¥ ACRIM = 01
¥ PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/027 - 2287.499274
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/046 show the observatory data loss to be 300 hours, 19
minutes, 59.848 seconds. This shows an overall increase of 6 minutes and 17.760
seconds since the last report period This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The increase of 6 minutes and 17.760 seconds occurred over three (3) days:
· 5 minutes and 14.608 seconds on Day 040 due to a late AOS,
· 37.472 seconds on Day 041 due to a Data Hit,
· 25.680 seconds on Day 042 due to a Data Hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS is using both the Fine Sun Sensor (FSS) and the Three Axis Magnetometer
(TAM) as inputs to UFLTR this week. The FOT is pleased with their performance,
but continues to make adjustments to the TAM alignment to improve its
performance and UARS's attitude.
MEETINGS
The FOT attended a UARS Anomaly Meeting on Tuesday, February 13, 2001.
The FOT attended an OIRB on Friday, February 16, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 15 March 2001 (DOY 074). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 February 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51496 042/0134 01-093 ACRIM Power ON
51497 042/0234 01-094 SUSIM "Power ON, Started"
51503 042/1154 01-095 PEM ZEPS/HEPS Day-Only
51506 042/1622 01-094 SUSIM Power ON Complete
51511 043/0053 01-096 AXIS "Power ON, 7w"
51521 043/1625 01-097 OBC MAGMOD Update
51536 044/1707 01-098 OBC TAM Update
51540 045/0001 01-099 HALOE "Power ON, part 1"
51543 045/0429 01-100 HALOE "Power ON, part 2"
51545 045/0755 01-101 NBTR NOOPS D-045
51552 045/1814 01-102 NBTR NOOPS D-046
51560 046/0652 01-102 NBTR NOOPS D-046
51566 046/1632 01-103 OBC TAM ALIGNMENT
51567 046/1816 01-104 NBTR NOOPS D-047
51571 047/0139 01-104 NBTR NOOPS D-047
51582 047/1857 01-105 NBTR NOOPS D-048
51585 047/2320 01-106 PEM/AXIS AXIS Power OFF
51583 047/2038 01-107 OBC Clear ERR flag
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 February 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 February 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 February 2001
09 Feb 2001 - GMT Day 040 Beta = 03.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ------ ------- -----
1 +000.0/-028.0 -3.68/-4.94 +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-61.6 +5.15/+3.87 +29.6/-13.6 24.8 1.020 21.0
3 +33.6/-50.4 +4.19/+2.92 +28.8/-15.6 24.8 1.010 20.7
16 Feb 2001 - GMT Day 047 Beta = 25.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ------ ------- -----
1 +000.0/-022.4 -3.05/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +22.4/-67.2 +5.47/+4.19 +29.6/-14.4 25.4 1.040 20.6
3 +28.0/-44.8 +4.51/+2.92 +29.2/-14.4 25.4 1.030 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 February 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-011 OBC 51581 047/1714Z N/A Incorrect ATC load.
No Impact.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-011 OBC 51581 047/1714Z N/A Incorrect ATC load.
No Impact.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-009 FOT see list Feb 01 14170 Generic Late Acquisition
SEVNTFW01/07